Ed and Dave are going to divide a cash prize of $14,000 in a ratio of 2 to 5. How much will Dave receive?
Found 2 solutions by adamchapman, amalm06:
Answer by adamchapman(301)

The money is beieng split into 2/7 and 5/7.
Dave gets (2/7)x14000=4000 dollars
I hope this helps,
Adam
P.S. please visit my website, it may be helpful to you. The address is www.geocities.com/quibowibbler

Answer by amalm06(224)

D/E=5/2, e.g. Dave receives 2.5 times the amount that Ed does.
Therefore, for every $7, Dave receives $5, or 5/7 of the total
If the total is $14000, Dave receives (5/7)(14000), or $10000 (Answer)